OpenStudy (anonymous):

Using math induction prove that (2n^3) less than or = (4^n). If N greater or equal 3

OpenStudy (freckles):

So you want to prove \[2n^3 \le 4n \text{ for } n \ge 3\]

OpenStudy (freckles):

show the base case first

OpenStudy (freckles):

that is the base case being n=3

OpenStudy (freckles):

it doesn't look like it holds for n=3

OpenStudy (freckles):

so you are done because it is false

OpenStudy (freckles):

any questions?

OpenStudy (freckles):

or is that not what you meant?

OpenStudy (anonymous):

i replaced n by 3 for 2n^3 gave me 54 and 4^3 gave me 81 soo in this case wouldn't it be true that the 4^n is greater for the case

OpenStudy (freckles):

So you are doing a different problem now?

OpenStudy (freckles):

I see 4n

OpenStudy (anonymous):

oh my mistake i meant for 4^n, sorry

OpenStudy (freckles):

\[2n^3 \le 4^n \text{ for } n \ge 3\]

OpenStudy (freckles):

so if this was the inequality should it is true for n=3 and yes this one would be true for the base case

OpenStudy (anonymous):

exactly i got stuck when replacing the n for (k+1)

OpenStudy (freckles):

now you suppose \[2k^3 \le 4^k \text{ for some integer k } \\ \text{ and we want to show } 2(k+1)^3 \le 4^{k+1}\]

OpenStudy (freckles):

hmm... now we think on different strategies

OpenStudy (anonymous):

i got to that part and then i'm stuck what to do next

OpenStudy (freckles):

just thinking \[2(k+1)^3 =2(k^3+3k^2+3k+1)\\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1)\]

OpenStudy (freckles):

\[2(k+1)^3 =2(k^3+3k^2+3k+1)\\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1) \\ \] well we know k>=3

OpenStudy (freckles):

thought i had an idea but it escaped me

OpenStudy (freckles):

just so we can think faster I will call @ganeshie8 here

OpenStudy (anonymous):

sounds good

OpenStudy (freckles):

\[2(k+1)^3 =2(k^3+3k^2+3k+1)\\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1) \\ \\ \le 4^k+2(kk^2+kk+1) =4^k+2(k^3+k^2+1)\]

OpenStudy (freckles):

it looks like we can use our induction hypothesis again

OpenStudy (freckles):

\[2(k+1)^3 =2(k^3+3k^2+3k+1)\\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1) \\ \\ \le 4^k+2(kk^2+kk+1) =4^k+2(k^3+k^2+1) \\ \le 4^k+4^k+2(k^2+1)=2 \cdot 4^k+2(k^2+1)\]

OpenStudy (freckles):

we want to show 2(k^2+1) is less than or equal to 2*4^k if we can

OpenStudy (freckles):

because 2*4^k+2*4^k=4*4^k=4^(k+1)

OpenStudy (freckles):

so we can simplify this is a little show now that k^2+1<=4^k somehow

OpenStudy (anonymous):

ok I see what you've done

OpenStudy (freckles):

i didn't finish it there is still a little something for you to do

ganeshie8 (ganeshie8):

\[\forall k \ge 1, ~~k^2+1 \le k^3 +k^3 = 2k^3 \le 4^k\]

OpenStudy (freckles):

yep

ganeshie8 (ganeshie8):

ah the last piece is by induction hypothesis so its not k>=1. it should be k>=3

OpenStudy (freckles):

I wonder if there is a shorter faster way

OpenStudy (freckles):

I guess I could have left that one thing as 2(3k+1) instead of wrote <=2(kk+1)=2(k^2+1)

OpenStudy (freckles):

I kinda like it as 3k<kk<k^3 though

OpenStudy (anonymous):

where would as 2(3k+1)? right before you factor 2(k+1)^3

OpenStudy (freckles):

\[2(k+1)^3 =2(k^3+3k^2+3k+1)\\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1) \\ \\ \le 4^k+2(kk^2+kk+1) =4^k+2(k^3+k^2+1) \\ \le 4^k+4^k+2(k^2+1)=2 \cdot 4^k+2(k^2+1) \] second line at the end the 2(3k^2+3k+1) or 2(3k^2)+2(3k+1)

OpenStudy (freckles):

your pieces are a bit scattered do you think you can piece it all together? does anything not make sense

OpenStudy (anonymous):

it's starting to make a lot more sense I just have to view all things that have been written so i don't get mixed up

OpenStudy (freckles):

i didn't know where i was going to go at first sometimes you just have to play with ideas i kinda think sometimes it helps to be a scatterbrain I could be wrong about it

OpenStudy (freckles):

it being the scatterbrain part

OpenStudy (freckles):

i usually call @ganeshie8 when I get nervous and scared :p

OpenStudy (anonymous):

no it's fine thank you very much , I needed help on this problem. the section where you wrote 4^k+2(3k^2...)you factored out the 3 and where did it go?

OpenStudy (freckles):

I will write a reason to the right or directly underneath each step

OpenStudy (freckles):

one sec

OpenStudy (freckles):

\[2(k+1)^3 =2(k^3+3k^2+3k+1) \text{ multiplied } \\ =2k^3+2(3k^2+3k+1) \le 4^k+2(3k^2+3k+1) \text{ use induction hypothesis } \\ \le 4^k+2(kk^2+kk+1) \text{ used } k \ge 3 \\ =4^k+2(k^3+k^2+1) \text{ did a little multipling } \\ \le 4^k+4^k+2(k^2+1) \text{ used induction hypothesis again } \\ =2\cdot 4^k+2(k^2+1) \text{ combined like terms } \\ \le 2 \cdot 4^k+2(k^3+k^3) \text{ \because } k^3<1 \text{ and } k^3>k^2 \text{ for } k \ge 3 \\ =2 \cdot 4^k+2(2k^3) \text{ combined like terms } \\ \le 2\cdot 4^k+2(4^k) \text{ used induction hypothesis }\]

OpenStudy (freckles):

last step is to basically combine like terms

OpenStudy (freckles):

and use law of exponents

OpenStudy (anonymous):

ohh wow thank you very much , this is perfect

OpenStudy (freckles):

you can correct the k^3>1 thing right

OpenStudy (freckles):

i put k^3<1

OpenStudy (anonymous):

ok will do thanks again

OpenStudy (freckles):

np